OpenStudy (anonymous):

f(x) = ln(\sqrt{{9x+4}/sqrt{4x-2}}) find f'(x)

OpenStudy (anonymous):

i got \[\frac{\frac{1}{2}\!\left(9x+4\right)^{-\frac{1}{2}}\cdot 9}{\sqrt{9x+4}}-\frac{\frac{1}{2}\!\left(4x-2\right)^{\frac{-1}{2}}\cdot 4}{\sqrt{4x-2}}\]

OpenStudy (anonymous):

but it says it's wrong

OpenStudy (anonymous):

\[f \left( x \right)=\frac{ 1 }{ 2 }\left\{ \ln \left( 9x+4 \right)-\ln \left( 4x-2 \right) \right\}\] \[f'\left( x \right)=\frac{ 1 }{2 }\left( \frac{ 9 }{9x+4 }-\frac{ 4 }{4x-2 } \right)\]

OpenStudy (anonymous):

what about \[f(x) = 4^x \log_{7} (x)\] will the answer be \[4^(x)*(1/\ln(7)*(1/x))\]

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

diff taking uv functions. d/dx(4^x)=4^x ln4

OpenStudy (anonymous):

so it would be 4^(x)ln4*(1/ln(7)*(1/x))

OpenStudy (anonymous):

d/dx(uv)=uv'+vu' d/dx log7 (x)=1/x *log7( e)

OpenStudy (anonymous):

thanks